OpenStudy (anonymous):

How do you rationalize the following denominator and reach a simplified solution? Please show detailed steps.

OpenStudy (anonymous):

\[(21-2\sqrt{21})\div(21+2\sqrt{21})\]

sam (.sam.):

Multiply with conjugate pairs

sam (.sam.):

(525-84root(21) ) / 357

sam (.sam.):

One of my rough calcualtions

OpenStudy (anonymous):

\[((21-2\sqrt{21})\times(21-2\sqrt{21})\div((21+2\sqrt{21})\times(21-2\sqrt{21}))\]\[(a-b)\times(a-b) = (a-b )^{2}\] \[(a-b)\times(a+b) = a ^{2}-b ^{2}\] \[(21-2\sqrt{21})^{2}\div(21^{2}-(2\sqrt{21})^{2})\] \[(21^{2}+(2\sqrt{21})^{2}-2\times21\times2\sqrt{21})\div(21^{2}-(2\sqrt{21})^{2})\] \[(441+4\times21-84\sqrt{21})\div(441-4\times21)\] \[(525-84\sqrt{21})\div(357)\] \[(25-4\sqrt{21})\div(17)\]

OpenStudy (anonymous):

Excellent, thank you very much.